# Graph Convolutions Enrich the Self-Attention in Transformers!

### v2

@@ Abstract
Transformers, renowned for their self-attention mechanism, have achieved state-of-the-art performance across various tasks in natural language processing, computer vision, time-series modeling, etc. However, one of the challenges with deep Transformer models is the oversmoothing problem, where representations across layers converge to indistinguishable values, leading to significant performance degradation. We interpret the original self-attention as a simple graph filter and redesign it from a graph signal processing (GSP) perspective. We propose graph-filter-based self-attention (GFSA) to learn a general yet effective one, whose complexity, however, is slightly larger than that of the original self-attention mechanism. We demonstrate that GFSA improves the performance of Transformers in various fields, including computer vision, natural language processing, graph pattern classification, speech recognition, and code classification.
@@ 1 Introduction
Figure 1: Our proposed GFSA performs better than selected Transformers in various domains. We achieve these results with only tens to hundreds of additional parameters to Transformers.
@@
Transformers are arguably one of the best feats in the field of deep learning. They are now showing state-of-the-art performance in various fields, ranging from computer vision to natural language processing to graph pattern classification to speech recognition, and so forth (Vaswani et al., 2017; Devlin et al., 2019; Radford et al., 2018; 2019; Dosovitskiy et al., 2021; Touvron et al., 2021a; Zhou et al., 2021a; Liu et al., 2021; Gulati et al., 2020; Latif et al., 2023; Ying et al., 2021; Rampášek et al., 2022). Recently, there have been several studies conducted on better understanding them (Gong et al., 2021; Ali et al., 2023; Wang et al., 2022); there exists a common agreement among researchers that the self-attention is one of the keys leading to the success.
@@
However, there also exist several studies pointing out potential limitations of the self-attention (Zhou et al., 2021a; Dong et al., 2021; Gong et al., 2021; Guo et al., 2023). For instance, Shi et al. (2022) revealed an analogy between the self-attention and the residual graph convolutional network (GCN), showing that BERT also suffers from a notorious problem of GCNs, called oversmoothing, i.e., tokens’ latent representations become similar to each other at the deeper layers. In every self-attention layer, value vectors are aggregated in a weighted average manner since each row-wise sum of the attention matrix is always 1. Although each self-attention layer has its own attention matrix, this aggregation method causes the oversmoothing problem, not only in Transformers but also in graph neural networks (Oono & Suzuki, 2020; Cai & Wang, 2020; Wang et al., 2021a; Rusch et al., 2023; Keriven, 2022; Zhou et al., 2021a; Gong et al., 2021; Yan et al., 2022; Noci et al., 2022; Shi et al., 2022; Wang et al., 2022; Ali et al., 2023; Wu et al., 2023a; b). However, we confine our discussion to the oversmoothing of Transformers (cf. Section 2).
Being inspired by them, we redesign the self-attention from the perspective of graph signal processing (GSP). However, performing graph convolutions in the self-attention layer may incur non-trivial computational overheads. Therefore, our key design point is to learn a general but effective graph filter with minimal overheads. In general, a graph filter on a graph $\mathcal{G}$ is represented by a polynomial expression based on its adjacency or Laplacian matrix — in this regard, the existing self-attention mechanism can be understood as the simplest graph filter with $\bar{\bm{A}}$ only, where $\bar{\bm{A}}\in[0,1]^{n\times n}$ means a learned attention matrix and $n$ is the number of input tokens.
Our proposed graph filter consists of an identity term and two matrix polynomial terms, $\bar{\bm{A}}$ and $\bar{\bm{A}}^{K}$ — one can design better filters with more polynomial terms but we avoid it since Transformers already require very large computation. The $K$-th power, $\bar{\bm{A}}^{K}$, may also require high computation when the number of tokens is large. To avoid this, we further approximate $\bar{\bm{A}}^{K}$ using the element-wise first-order Taylor approximation. Therefore, one can consider that our proposed graph filter is the very next complicated filter after the one used by the original self-attention mechanism. However, its efficacy is tremendous for various Transformers in various fields (cf. Figure 1).
Our proposed filter enriches the self-attention with more diverse frequency information (see Figure 2(a)) — low (resp. high) frequency signals on $\mathcal{G}$ mean neighboring nodes have similar (resp. different) values. Therefore, our method is able to not only effectively address the oversmoothing problem but also learn better latent representations for downstream tasks.
There exist a couple of prior works to enrich the self-attention mechanism with high frequency information (Wang et al., 2022; Bai et al., 2022). In comparison with them, our proposed graph filter is distinctive in the following aspects: i) our proposed filter is more effective and shows better performance with comparable computational overheads, ii) our proposed filter is well-aligned with recent advancements in the GCN community — in other words, some graph filters used by recent advanced GCN methods are special cases of our proposed graph filter, which is not the case for prior works, and iii) other methods were typically studied for certain domains only whereas we test our method in 6 domains — for instance, DiversePatch (Gong et al., 2021) works only for Vision Transformers.
We replace the self-attention layer of selected Transformers in various fields with our proposed graph filter-based layer without changing other parts. Therefore, the accuracy increases in them are solely by our proposed graph filter-based self-attention. These enriched Transformers increase the model performance by 1.63% for image classification, 6.25% for natural language understanding, 0.31% for causal language modeling, 4.03% for graph classification, 4.76% for speech recognition, and 2.40% for code classification (see Figure 1). We summarize our core contributions as follows:
  • We propose a graph filter-based self-attention (GFSA) mechanism that is general yet effective. GFSA learns a graph filter with an identity term and two polynomial terms, which is more effective than the simple self-attention mechanism.
  • We demonstrate that GFSA can significantly improve the performance of Transformers on a variety of tasks. GFSA achieves improved results on computer vision, natural language processing, graph pattern classification, speech recognition, and code classification.
2 Background & Related Work
2.1 Self-Attention in Transformers
The core building block of the Transformer architecture is the self-attention mechanism, which enables the model to learn attention patterns over its input tokens (Vaswani et al., 2017). The self-attention mechanism, denoted as $\textrm{SA}:\mathbb{R}^{n\times d}\rightarrow\mathbb{R}^{n\times d}$, can be expressed as follows:
$\textrm{SA}(\bm{X})=\mathrm{softmax}\Big{(}\frac{\bm{X}\bm{W}_{\textrm{key}}(\bm{X}\bm{W}_{\textrm{qry}})^{\intercal}}{\sqrt{d}}\Big{)}\bm{X}\bm{W}_{\textrm{val}}=\bar{\bm{A}}\bm{X}\bm{W}_{\textrm{val}},$ (1)
2.2 Self-Attention and Graph Convolutional Filter
The self-attention matrix used in Transformers can be seen as a symmetrically normalized adjacency matrix of a graph (Shi et al., 2022; Guo et al., 2023; Maskey et al., 2023). A weighted graph $\mathcal{G}$ with adjacency matrix $\bm{A}$ can be constructed by using the input tokens as $n$ nodes and the edge weights between node $i$ and node $j$ as $\exp((\bm{X}\bm{W}_{\textrm{qry}})^{\intercal}(\bm{X}\bm{W}_{\textrm{key}}))$. We can rewrite the self-attention matrix $\bar{\bm{A}}$ as $\frac{\exp{(\bm{X}\bm{W}_{\textrm{qry}})^{\intercal}_{i}(\bm{X}\bm{W}_{\textrm{key}})_{j}}}{\sum_{k=1}^{d}\exp{(\bm{X}\bm{W}_{\textrm{qry}})^{\intercal}_{i}(\bm{X}\bm{W}_{\textrm{key}})_{k}}}$. This allows $\bar{\bm{A}}$ to be interpreted as the symmetrically normalized adjacency matrix. In other words, $\bar{\bm{A}}=\bm{D}^{-1}\bm{A}$, where $\bm{D}=\textrm{diag}(d_{1},d_{2},\ldots,d_{n})$ and $d_{i}=\sum_{j}\bm{A}_{i,j}$.
Our new attention method is designed on top of graph signal processing (GSP) which has a close connection to discrete signal processing (DSP) (Sandryhaila & Moura, 2013; 2014). In DSP, a discrete signal with a length of $n$ can be represented by a vector $\bm{x}\in\mathbb{R}^{n}$. Let $\bm{g}\in\mathbb{R}^{n}$ be a filter that we want to apply to $\bm{x}$. The convolution $\bm{x}*\bm{g}$ can be written as follows:
$\bm{y}_{i}=\sum_{j=1}^{n}\bm{x}_{j}\bm{g}_{i-j},$ (2)
GSP can be understood as a generalized concept of DSP. Signals are defined on the nodes of a graph, and the graph’s structure influences signal processing operations. In addition, the graph convolution filter with $n$ nodes can be written with a shift operator $\bm{S}$ as follows:
$\bm{y}=\sum_{k=0}^{K}w_{k}\bm{S}^{k}\bm{x}=\bm{H}\bm{x},$ (3)
where $\bm{x}\in\mathbb{R}^{n}$ is a 1-dimensional graph signal, $K$ is the maximum order of polynomial, and $w_{k}\in[-\infty,\infty]$ is a coefficient. $\bm{S}$ is an $n\times n$ matrix where $(i,j)$-th element is non-zero if and only if there is an edge from node $i$ to $j$. Two representative samples of $\bm{S}$ are adjacency and Laplacian matrices. The linear and shift-invariant graph filter $\bm{H}$ is the same as $\sum_{k=0}^{K}w_{k}\bm{S}^{k}$ with a large enough value of $K$, which is called matrix polynomial (Marques et al., 2020). We note that this graph filtering operation can be extended to $d$-dimensional cases as in  Equation 1. Being inspired by (Zou et al., 2022; Maskey et al., 2023), we rely on the singular value domain analysis to understand the low/high-pass filter characteristics of filters (cf. Fig. 2). See more discussion in Appendix N.
In the context of self-attention within Transformers, the core part of the self-attention in Equation 1, i.e., $\bar{\bm{A}}\bm{X}$, can be considered as a $d$-dimensional graph filter with $\bar{\bm{A}}$ only, where $\bm{H}=\bar{\bm{A}}$. Our goal in this paper is design a simple (for computational purposes) yet effective form of $\bm{H}$.
2.3 Oversmoothing in GCNs and Transformers
Oversmoothing is a phenomenon observed in deep learning models, especially in GCNs (Kipf & Welling, 2017; Veličković et al., 2018). As information is aggregated over multiple layers for multiple nodes (tokens), latent representations tend to become similar to each other, leading to a loss of distinctiveness in the representations (Oono & Suzuki, 2020; Zhou et al., 2020; Rusch et al., 2023).
Surprisingly, an oversmoothing-like phenomenon is also observed in Transformers (Wang et al., 2022; Shi et al., 2022). Unlike CNNs, Transformers can not benefit from simply deepening layers after a certain depth. Earlier studies hypothesize that this may be due to issues such as the attention or feature collapse or due to uniformity among patches or tokens (Zhou et al., 2021a; Gong et al., 2021; Yan et al., 2022). Dong et al. (2021) also point out that the output of a pure Transformer, i.e., an attention mechanism without skip connections or MLPs, tends to converge to a rank-1 matrix (Dong et al., 2021). This analysis is followed by (Noci et al., 2022), which suggests that rank collapses incur vanishing gradients of attention queries and keys.
In this context, the self-attention acts as a low-pass filter, since the self attention calculates the weighted average of the value vectors of tokens. Wang et al. (2022, Theorem 1) also reveal that the self-attention is a low-pass filter, continuously reducing high-frequency information. This nature contributes to the oversmoothing phenomenon as unique high-frequency features are lost in deeper layers of the network, further worsening the uniformity of token representations. Therefore, we extend the term “oversmoothing” to describe the degeneration challenge observed in Transformers.
There have been proposed many empirical countermeasures for Vision Transformer (ViT), such as patch diversification (Zhou et al., 2021b; Gong et al., 2021), rank collapse alleviation (Zhou et al., 2021a; Zhang et al., 2021), and training stabilization (Touvron et al., 2021a; Zhai et al., 2023). Similar alleviating methods have been also proposed in the field of NLP, such as unsupervised learning (Chen et al., 2023), and resolve the oversmoothing and the token uniformity (or information diffusion) problems (Dong et al., 2021; Yan et al., 2022). There are studies on utilizing high frequency information via frequency domain analyses (Wang et al., 2022; Bai et al., 2022), but they are not designed on top of graph filtering perspectives. This paper addresses the oversmoothing problem with graph filters since the self-attention mechanism is a basic graph filtering operation as seen in the previous subsection.
3 Proposed Method
@@
Figure 2: Filter frequency response, cosine similarity, and singular values on ImageNet-1k for DeiT-S and DeiT-S + GFSA. Details and more visualizations are in Appendix D.
3.1 Graph Filter-based Self-Attention Layers
@@
Let $\bar{\bm{A}}\in[0,1]^{n\times n}$, where $n$ is the number of tokens in the input to the self-attention layer, be a self-attention matrix. Since transformers use multi-head self-attentions, there are multiple such matrices. For simplicity but without loss of generality, we discuss only one head in one layer.
In the perspective of GSP, $\bar{\bm{A}}$ corresponds to a special case where i) nodes are completely connected, and ii) $\bar{\bm{A}}$ is normalized. $\bar{\bm{A}}$ can be used as a shift operator. When $K$ is large, however, it is hard to calculate $\bar{\bm{A}}^{K}$. We use the following element-wise first-order Taylor approximation:
$\bar{\bm{A}}^{K}\approx\bar{\bm{A}}+(K-1)(\bar{\bm{A}}^{2}-\bar{\bm{A}}),$ (4)
@@
where $\bar{\bm{A}}^{2}-\bar{\bm{A}}$ is an approximated derivative at the position of $K=1$ (see details in Appendix P). At the end, we propose to use the following graph filter where the two lowest-order terms and one high-order term of the matrix polynomial are used:
$\bm{H}:=w_{0}\bm{I}+w_{1}\bar{\bm{A}}+w_{K}(\bar{\bm{A}}+(K-1)(\bar{\bm{A}}^{2}-\bar{\bm{A}})),$ (5)
@@
where $w_{0}$, $w_{1}$, $w_{K}$ are coefficients and $K$ is a hyper-parameter. The coefficients can be learnable weights and we learn them with gradient descent algorithms.
@@
Our proposed graph filter-based self-attention (GFSA) is defined with the graph filter $\bm{H}$ as follows:
$\textrm{GFSA}(\bm{X})=\bm{H}\bm{X}\bm{W}_{\textrm{val}}.$ (6)
@@
We replace the original self-attention layer in various Transformers with the proposed graph filter-based layer without changing other parts. Therefore, GFSA can be plugged into any Transformers that rely on the self-attention.
3.2 Theoretical Analysis

### v5

@@ Abstract
Transformers, renowned for their self-attention mechanism, have achieved state-of-the-art performance across various tasks in natural language processing, computer vision, time-series modeling, etc. However, one of the challenges with deep Transformer models is the oversmoothing problem, where representations across layers converge to indistinguishable values, leading to significant performance degradation. We interpret the original self-attention as a simple graph filter and redesign it from a graph signal processing (GSP) perspective. We propose a graph-filter-based self-attention (GFSA) [^1] The source code of GFSA is available at: https://github.com/jeongwhanchoi/GFSA. to learn a general yet effective one, whose complexity, however, is slightly larger than that of the original self-attention mechanism. We demonstrate that GFSA improves the performance of Transformers in various fields, including computer vision, natural language processing, graph-level tasks, speech recognition, and code classification.
@@ 1 Introduction
Figure 1: Performance improvements (%) of our GFSA when integrated with different Transformer backbones in various domains. We achieve these results with only tens to hundreds of additional parameters to Transformers.
@@
Transformers are arguably one of the best feats in the field of deep learning. They are now showing state-of-the-art performance in various fields, ranging from computer vision to natural language processing, prediction tasks on graphs, speech recognition, and so forth [77, 16, 60, 61, 19, 74, 101, 45, 27, 41, 63, 51, 59, 38, 95, 72]. Recently, there have been several studies conducted on better understanding them [25, 3, 80]; there exists a common agreement among researchers that the self-attention is one of the keys leading to the success.
@@
However, there also exist several studies pointing out potential limitations of the self-attention [101, 18, 25, 28]. For instance, Shi et al. [71] revealed an analogy between the self-attention and the residual graph convolutional network (GCN), showing that BERT also suffers from a notorious problem of GCNs, called oversmoothing, i.e., tokens’ latent representations become similar to each other at the deeper layers. In every self-attention layer, value vectors are aggregated in a weighted average manner since each row-wise sum of the attention matrix is always 1. Although each self-attention layer has its own attention matrix, this aggregation method causes the oversmoothing problem, not only in Transformers but also in graph neural networks [54, 7, 79, 66, 37, 101, 25, 92, 53, 71, 80, 3, 89, 90]. However, we confine our discussion to the oversmoothing of Transformers (see Section 2).
@@
Figure 2: Filter frequency response, cosine similarity, and singular values on ImageNet-1k for DeiT-S and DeiT-S + GFSA. Details and more visualizations are in Appendices C and D.
Being inspired by them, we redesign the self-attention from the perspective of graph signal processing (GSP) — in particular, we resort to GSP on directed graphs since the attention matrix is asymmetric. However, performing graph convolutions in the self-attention layer may incur non-trivial computational overheads. Therefore, our key design point is to learn a general but effective graph filter with minimal overhead. In general, a graph filter on a graph $\mathcal{G}$ is represented by a polynomial expression based on its adjacency or Laplacian matrix — in this regard, the existing self-attention mechanism can be understood as the simplest graph filter with $\bar{\bm{A}}$ only, where $\bar{\bm{A}}\in[0,1]^{n\times n}$ means a learned attention matrix and $n$ is the number of input tokens.
Our proposed graph filter consists of an identity term and two matrix polynomial terms, $\bar{\bm{A}}$ and $\bar{\bm{A}}^{K}$. One can design better filters with more polynomial terms, but we avoid it since Transformers already require very large computation. The $K$-th power, $\bar{\bm{A}}^{K}$, may also require a high computation when the number of tokens is large. To avoid this, we further approximate $\bar{\bm{A}}^{K}$ using the element-wise first-order Taylor approximation. Therefore, one can consider that our proposed graph filter is the very next complicated filter after the one used by the original self-attention mechanism. However, its efficacy is tremendous in various fields (cf. Fig. 1).
Our proposed filter enriches the self-attention with more diverse frequency information (see Fig. 2(a)) — low (resp. high) frequency signals on $\mathcal{G}$ mean neighboring nodes have similar (resp. different) values. Therefore, our method is able to not only effectively address the oversmoothing problem but also learn better latent representations for downstream tasks.
There exist a couple of prior works to enrich the self-attention mechanism with high frequency information [80, 4]. In comparison with them, our proposed graph filter is distinctive in the following aspects: i) our proposed filter is more effective and shows better performance with comparable computational overheads, ii) our proposed filter is well-aligned with recent advancements in the GCN community — in other words, some graph filters used by recent advanced GCN methods are special cases of our proposed graph filter, which is not the case for prior works, and iii) other methods were typically studied for certain domains only whereas we test our method in 6 domains — for instance, DiversePatch [25] works only for Vision Transformers (ViTs).
We replace the self-attention layer of selected Transformers in various fields with our proposed graph filter-based layer without changing other parts. Therefore, the accuracy increases in them are solely by our proposed graph filter-based self-attention. These enriched Transformers increase the model performance by 1.63% for image classification, 6.25% for natural language understanding, 0.31% for causal language modeling, 4.03% for graph regression, 4.76% for speech recognition, and 2.40% for code classification (see Fig. 1). Our core contributions are as follows:
  • We propose a graph filter-based self-attention (GFSA) mechanism, integrating an identity term and two polynomial terms for general yet effective than the simple self-attention mechanism (Section 3).
  • We demonstrate that GFSA improves the performance of Transformers on a variety of tasks. GFSA achieves improved results on natural language processing, computer vision, speech recognition, graph-level tasks, and code classification (Sections 5.1 to 5.6).
  • We devise a strategy to selectively apply GFSA to even-numbered layers, effectively mitigating the computational overhead while preserving GFSA’s performance (Section 6).
2 Background & Related Work
2.1 Self-Attention in Transformers
The core building block of the Transformer architecture is the self-attention mechanism, which enables the model to learn attention patterns over its input tokens [77]. The self-attention mechanism, denoted as $\textrm{SA}:\mathbb{R}^{n\times d}\rightarrow\mathbb{R}^{n\times d}$, can be expressed as follows:
$\textrm{SA}(\bm{X})=\textrm{softmax}\Big{(}\frac{\bm{X}\bm{W}_{\textrm{key}}(\bm{X}\bm{W}_{\textrm{qry}})^{\intercal}}{\sqrt{d}}\Big{)}\bm{X}\bm{W}_{\textrm{val}}=\bar{\bm{A}}\bm{X}\bm{W}_{\textrm{val}},$ (1)
2.2 Self-Attention and Graph Convolutional Filter
The self-attention matrix used in Transformers has the form of symmetrically normalized adjacency matrix where each token become a node [71, 28] — the symmetrically normalized adjacency matrix is a special case of asymmetric (or directed) adjacency matrix where each row is normalized and is frequently used for the graph signal processing (GSP) on directed graphs [49]. A weighted graph $\mathcal{G}$ with adjacency matrix $\bm{A}$ can be constructed by using the input tokens as $n$ nodes and the edge weights between node $i$ and node $j$ as $\exp((\bm{X}\bm{W}_{\textrm{qry}})^{\intercal}(\bm{X}\bm{W}_{\textrm{key}}))$. We can rewrite the self-attention matrix $\bar{\bm{A}}$ as $\frac{\exp{((\bm{X}\bm{W}_{\textrm{qry}})^{\intercal}_{i}(\bm{X}\bm{W}_{\textrm{key}})_{j}})}{\sum_{k=1}^{d}\exp{(\bm{X}\bm{W}_{\textrm{qry}})^{\intercal}_{i}(\bm{X}\bm{W}_{\textrm{key}})_{k}}}$. This allows $\bar{\bm{A}}$ to be interpreted as the symmetrically normalized adjacency matrix. In other words, $\bar{\bm{A}}=\bm{D}^{-1}\bm{A}$, where $\bm{D}=\textrm{diag}(d_{1},d_{2},\ldots,d_{n})$ and $d_{i}=\sum_{j}\bm{A}_{i,j}$.
Our new attention method is designed on top of GSP which has a close connection to discrete signal processing (DSP) [67, 68]. In DSP, a discrete signal with a length of $n$ can be represented by a vector $\bm{x}\in\mathbb{R}^{n}$. Let $\bm{g}\in\mathbb{R}^{n}$ be a filter that we want to apply to $\bm{x}$. The convolution $\bm{x}*\bm{g}$ can be written as follows:
$\bm{y}_{i}=\sum_{j=1}^{n}\bm{x}_{j}\bm{g}_{i-j},$ (2)
GSP can be understood as a generalized concept of DSP. Signals are defined on the nodes of a graph, and the graph’s structure influences signal processing operations. In addition, the linear and shift-invariant graph convolution filter $\bm{H}$ with $n$ nodes can be written with a shift operator $\bm{S}$ as follows — $\bm{S}$ can be from a directed graph [48]:
$\bm{y}=\bm{H}\bm{x}=\sum_{k=0}^{K}w_{k}\bm{S}^{k}\bm{x},$ (3)
where $\bm{x}\in\mathbb{R}^{n}$ is a 1-dimensional graph signal, $K$ is the maximum order of polynomial, and $w_{k}\in[-\infty,\infty]$ is a coefficient. $\bm{S}$ is an $n\times n$ matrix where $(i,j)$-th element is non-zero if and only if there is an edge from node $i$ to $j$. Two representative samples of $\bm{S}$ are adjacency and Laplacian matrices. The graph filter $\bm{H}$ is the same as $\sum_{k=0}^{K}w_{k}\bm{S}^{k}$ with a large enough value of $K$, which is called matrix polynomial [48]. We note that this graph filtering operation can be extended to $d$-dimensional cases as in Eq. (1). Being inspired by Zou et al. [106] and Maskey et al. [49], we rely on the singular value domain analysis to understand the low/high-pass characteristics of filters on directed graphs (cf. Fig. 2). See more discussion in Appendices E and F.
In the context of the self-attention within Transformers, the core part of the self-attention in Eq. (1), i.e., $\bar{\bm{A}}\bm{X}$, can be considered as a $d$-dimensional graph filter with $\bar{\bm{A}}$ only, where $\bm{H}=\bar{\bm{A}}$. Our goal in this paper is to design a simple (for computational purposes) yet effective form of $\bm{H}$.
2.3 Oversmoothing in GCNs and Transformers
Oversmoothing is a phenomenon observed in deep learning models, especially in GCNs [39, 78]. As information is aggregated over multiple layers for multiple nodes (tokens), latent representations tend to become similar to each other, leading to a loss of distinctiveness in the representations [54, 104, 66].
Surprisingly, an oversmoothing-like phenomenon is also observed in Transformers [80, 71]. Unlike CNNs, Transformers can not benefit from simply deepening layers after a certain depth. Earlier studies hypothesize that this may be due to issues such as the attention or feature collapse or due to uniformity among patches or tokens [101, 25, 92]. Dong et al. [18] also point out that the output of a pure Transformer, i.e., an attention mechanism without skip connections or MLPs, tends to converge to a rank-1 matrix [18]. This analysis is followed by [53], which suggests that rank collapses incur vanishing gradients of attention queries and keys.
In this context, the self-attention acts as a low-pass filter, since the self-attention calculates the weighted average of the value vectors of tokens. Wang et al. [80, Theorem 1] also reveal that the self-attention is a low-pass filter, continuously reducing high-frequency information. This nature contributes to the oversmoothing phenomenon as unique high-frequency features are lost in deeper layers of the network, further worsening the uniformity of token representations. Therefore, we extend the term “oversmoothing” to describe the degeneration challenge observed in Transformers.
There have been proposed many empirical countermeasures for ViT, such as patch diversification [102, 25], rank collapse alleviation [101, 99], and training stabilization [74, 98]. Similar alleviating methods have been also proposed in the field of NLP, such as unsupervised learning [9], and resolve the oversmoothing and the token uniformity (or information diffusion) problems [18, 92]. There are studies on utilizing high frequency information via frequency domain analyses [80, 4], but they are not designed on top of graph filtering perspectives. Dovonon et al. [20] find that Transformers are not inherently low-pass filters, but oversmoothing depends on the eigenspectrum of the self-attention layers. They propose a reparametrization of the Transformer weights, ensuring that oversmoothinng does not occur.
Our paper addresses the oversmoothing problem with graph filters since the self-attention mechanism is a basic graph filtering operation as seen in the previous subsection.
3 Graph Filter-based Self-Attention Layers
@@
Let $\bar{\bm{A}}\in[0,1]^{n\times n}$, where $n$ is the number of tokens in the input to the self-attention layer, be a self-attention matrix. Since Transformers use multi-head self-attentions, there are multiple such matrices. For simplicity, but without loss of generality, we discuss only one head in one layer.
From the GSP perspective, using $\bar{\bm{A}}$ as the shift operator, a graph filter can be represented as a matrix polynomial filter, as mentioned in Section 2.2. We aim to design this matrix polynomial filter using the two lowest-order terms and one high-order term in Eq. (3). The following theorem shows that, despite using the three terms, the filter can be either a low-pass filter or a high-pass filter, depending on the coefficient values.
Theorem 3.1 (Filter characteristics based on coefficient values).
Let $\bar{\bm{A}}$ be a self-attention matrix interpreted as a graph with connected components. Consider the polynomial graph filter defined by $\sum_{k=0}^{K}w_{k}\bar{\bm{A}}^{k}$, where $w_{2},w_{3},\ldots,w_{K-1}=0$ and only $w_{0}$, $w_{1}$, and $w_{K}$ are non-zero. If the coefficients $w_{k}$ for $k=0,1,K$ are positive and their sum is 1, then the polynomial filter acts as a low-pass filter, attenuating high-frequency components and promoting smoothness across the graph. Conversely, if $w_{k}=(-\alpha)^{k}$ for $k=0,1,K$ and $\alpha\in(0,1)$ with sufficient large $K$, the polynomial filter exhibits high-pass filter behavior.
@@
The proof of Theorem 3.1 is in Appendix G. Based on Theorem 3.1, we propose to use the following graph filter, $\bm{H}_{\text{GFSA}}$, where the two lowest-order terms and one high-order term of the matrix polynomial are used:
$\bm{H}_{\text{GFSA}}=w_{0}\bm{I}+w_{1}\bar{\bm{A}}+w_{K}\bar{\bm{A}}^{K},$ (4)
@@
where $w_{0}$, $w_{1}$, $w_{K}$ are coefficients and $K$ is a hyper-parameter where $K\geq 2$. The coefficients can be learnable weights and we learn them with gradient descent algorithms.
Approximation of the high-order term.
In Eq. (4), it is costly to calculate $\bar{\bm{A}}^{K}$ when $K$ is large, so we need a way to approximate the high-order term $\bar{\bm{A}}^{K}$ in GFSA. We use the first-order Taylor approximation at point $a=1$ for this purpose:
$f(x)\simeq f(a)+f^{\prime}(a)(x-a),$ (5)
thus, we approximate $f(K)=\bar{\bm{A}}^{K}$ as follows:
$f(K)=\bar{\bm{A}}^{K}\simeq f(1)+f^{\prime}(1)(K-1).$ (6)
Computing the derivative of $\bar{\bm{A}}^{K}$ directly at the evaluation point requires high computational costs. To overcome this problem, we adopt the forward finite difference method, which approximates derivatives with the difference term:
$f^{\prime}(K)=\frac{f(K+h)-f(K)}{h}=\frac{\bm{A}^{K+h}-\bm{A}^{K}}{h},$ (7)
where the approximation error is $\mathcal{O}(h^{2})$. To balance the trade-off between computational efficiency [^2] Calculating the power of a matrix for small $h$ requires a high computational cost since it is calculated by diagonalizing the matrix or using Schur normal form [31]. and accuracy, we set $h=1$. This method is inspired by the approach in  Brouwer et al. [6], which uses the difference term between two consecutive hidden states in discrete time to approximate the derivatives. Therefore, we approximate $\bar{\bm{A}}^{K}$ as:
$\begin{split}f(K)=\bar{\bm{A}}^{K}&\simeq f(1)+(\bar{\bm{A}}^{2}-\bar{\bm{A}})(K-1)\\ &=\bar{\bm{A}}+(K-1)(\bar{\bm{A}}^{2}-\bar{\bm{A}}).\end{split}$ (8)
The approximation for $\bar{\bm{A}}^{K}$ with $\bar{\bm{A}}$ and $\bar{\bm{A}}^{2}$ provides a simpler computation that can significantly reduce the required computational resources and time.
GFSA: our graph filter-based self-attention.
@@
Our proposed graph filter-based self-attention (GFSA) is defined with the graph filter $\tilde{\bm{H}}_{\text{GFSA}}$ as follows:
$\textrm{GFSA}(\bm{X})$ $:=\tilde{\bm{H}}_{\text{GFSA}}\bm{X}\bm{W}_{\textrm{val}},$ (9) $\tilde{\bm{H}}_{\text{GFSA}}$ $=w_{0}\bm{I}+w_{1}\bar{\bm{A}}+w_{K}(\bar{\bm{A}}+(K-1)(\bar{\bm{A}}^{2}-\bar{\bm{A}})),$ (10)
@@
where the last term is the approximated $\bar{\bm{A}}^{K}$ from Eq. (8). We replace the original self-attention layer in various Transformers with the proposed graph filter-based layer without changing other parts. Therefore, GFSA can be plugged into any Transformers that rely on the self-attention. For pseudocode, see Appendix I.
4 Properties of GFSA
This section analyzes the theoretical error of the $\bar{\bm{A}}^{K}$ approximation used by GFSA and how GFSA can mitigate oversmoothing. We also explain the meaning of GFSA’s high-order term in the context of Transformers and provide comparisons of GFSA in other models.
Theoretical characteristics of approximation error in GFSA.